OpenStudy (anonymous):

negative the square root of nine equals negative nine negative three 3 9

OpenStudy (anonymous):

\[\sqrt{-9}\]

OpenStudy (anonymous):

\[i = \sqrt{-1}\]

OpenStudy (anonymous):

so its 9?

OpenStudy (anonymous):

\[= 3i\]

OpenStudy (anonymous):

\[\sqrt{-9} = \sqrt{-1} * \sqrt{9}\] \[\sqrt{9} = 3\]

OpenStudy (anonymous):

so its 3?

OpenStudy (anonymous):

\[-\sqrt{9}=(-1)*\sqrt{9}=(-1)*3=-3 and \it is also equal \to (-1)*(-3)=3\]

OpenStudy (anonymous):

most of the time people just want the postive number if you take the squareroot - so if only one answer is correct I would take -3

OpenStudy (anonymous):

\[-\sqrt{9}\ or\ \sqrt{-9}?\]

OpenStudy (anonymous):

@yoyogators which is it?

OpenStudy (anonymous):

Okay so the answer is 3

OpenStudy (anonymous):

\[-\sqrt{9}\]

OpenStudy (anonymous):

Depending which of the two square roots you are referring to... if its\[\sqrt{-9}\]then yes.

OpenStudy (anonymous):

If it's \[-\sqrt{9} \] then it's like saying \[-1 \times \sqrt{9}\]

OpenStudy (anonymous):

\[\sqrt{9} = 3 \]3 * -1 = negative 3